OpenStudy (chrisneedshelp):

1/4 + 5/12 =? FRACTION HELP ANSWER THEN SIMPLIFY

OpenStudy (lgbasallote):

hint: \[\large \frac 14 \implies \frac{3}{12}\]

OpenStudy (chrisneedshelp):

Convert these unlike fractions to equivalent like fractions and add them. You must use the LCD to get the answer correct. If possible, reduce the final sum.

OpenStudy (lgbasallote):

psst look at the post above yours :S lol

OpenStudy (chrisneedshelp):

Convert these unlike fractions to equivalent like fractions and add them. You must use the LCD to get the answer correct. If possible, reduce the final sum. 1/4= ? 5/12=?

OpenStudy (chrisneedshelp):

LOL

OpenStudy (lgbasallote):

i think i just gave a key hint to this problem ;)

OpenStudy (chrisneedshelp):

lol i need more tho, :P

OpenStudy (chrisneedshelp):

though*

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (chrisneedshelp):

like, it says Convert these unlike fractions to equivalent like fractions and add them. You must use the LCD to get the answer correct. If possible, reduce the final sum.

OpenStudy (chrisneedshelp):

1/4=?

OpenStudy (chrisneedshelp):

and 5/12=? lol

OpenStudy (lgbasallote):

hmm did you read my first post?

OpenStudy (chrisneedshelp):

yes thanks for that lol :) but 5/12=?

OpenStudy (lgbasallote):

what is the denominator of 5/12?

OpenStudy (chrisneedshelp):

12

OpenStudy (lgbasallote):

is that the same as the first one?

OpenStudy (chrisneedshelp):

hmmmmm

OpenStudy (chrisneedshelp):

challenging question... no lol as in 1/4?

OpenStudy (lgbasallote):

by "first one" i mean the denominator of the fraction i wrote a while ago

OpenStudy (chrisneedshelp):

ohhhhhh

OpenStudy (chrisneedshelp):

yes

OpenStudy (chrisneedshelp):

3/12

OpenStudy (lgbasallote):

then no more need to change. it stays

OpenStudy (lgbasallote):

so now..since they are now common denominators you can add them 3/12 + 5/12 just copy the denominator then add the numerators

OpenStudy (chrisneedshelp):

:D)

OpenStudy (chrisneedshelp):

A LIFESAVER!!!

OpenStudy (lgbasallote):

glad to have hlped ^_^

OpenStudy (chrisneedshelp):

thanks :) i have more lol try to help im going to postthem up

OpenStudy (anonymous):

(1/5=3/12 ? 2/3=5/12)+8/24????